# WhiSPA: Semantically and Psychologically Aligned Whisper with Self-Supervised Contrastive and Student-Teacher Learning

### v1

@@ 5 Results & Discussion
We consider two popular speech models as baselines, Wav2Vec 2.0 Baevski et al. (2020) and whisper-tiny Radford et al. (2022), which are referred to as Wav2Vec-768 and Whisper-384 respectively. We measured the effectiveness of these embeddings by computing Pearson correlation coefficient ($r$) and mean squared error ($mse$) over a 10-fold cross-validated ridge regression model for each task. For each model variant in Table 2, we encode audio segments for each participant and aggregate them with a statistical mean to represent person-level embeddings for tasks in Table 3.
Audio Model WTC HiTOP PCL REX AVO NAM HYP INT DIS ANT SOM THD DET $r(\uparrow)$ $mse(\downarrow)$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ Wav2Vec-768 .23 128.44 .19 11.87 .15 3.93 .21 10.68 .20 16.62 .37 .20 .36 .23 .22 .12 -.02 .24 .21 .11 .11 .22 Whisper-384 .23 128.85 .21 11.77 .06 4.00 .19 10.87 .23 16.41 .39 .19 .33 .24 .33 .11 .07 .23 .28 .11 .29 .20 WhiSA-384 .29$\dagger$ 119.68 .27$\dagger$ 11.26 .19$\dagger$ 3.90 .26$\dagger$ 10.12 .28$\dagger$ 15.56 .55$\dagger$ .16 .53$\dagger$ .19 .43$\dagger$ .10 .22$\dagger$ .23 .37$\dagger$ .10 .33$\dagger$ .18 WhiSPA-384r .34$\dagger$ 119.24 .30$\dagger$ 11.23 .17 3.88 .31$\dagger$ 10.08 .32$\dagger$ 15.54 .56$\dagger$ .15 .53$\dagger$ .19 .42$\dagger$ .10 .23* .22 .39$\dagger$ .10 .39$\dagger$ .19 WhiSPA-394 .35$\dagger$ 118.91 .30$\dagger$ 11.18 .20 3.85 .32$\dagger$ 10.09 .32$\dagger$ 15.48 .57$\dagger$ .15 .54$\dagger$ .19 .43$\dagger$ .10 .$\dagger .2$37$\dagger$ .10 .38$\dagger$ .19
[…]
@@
Figure 3: Bivariate KDE contour plot of dimensionally reduced speech/text embeddings with PCA. Speech representations in blue. Text representations in red.
Between semantic and semantic-psychological alignment, the latter offered marginally better performance than the former, increasing by 1-5 Pearson points on both datasets. Interestingly, deriving psychological estimates from semantic dimensions (WhiSPA-394) was consistently better than the replacement (WhiSPA-384r) of 10 semantic dimensions with PsychEmb dimension. This shows the importance of curating the semantic dimensions before replacing them with different embeddings.
@@
Semantic-Psychological alignment is SotA for person-level psychological assessments.
Table 3 shows that the improvements brought by our aligned models over traditional models were preserved even when evaluated on a spectrum of downstream psychological assessment tasks. In particular, the alignment showed a stark increase in capturing deeper psychological conditions such as Internalizing ($\geq 16$ Pearson points) and Disinhibition ($\geq 20$ Pearson points) from very long durations of speech data. Consistent with behaviours exhibited with PsychEmb dimensions, in Table 2, semantic-psychological alignment from semantically-derived psychological dimensions (WhiSPA-394) performed the best, followed by semantic-psychological alignment from replacement (WhiSPA-384r) and finally semantic-only alignment (WhiSA-384). For these tasks, we averaged the segment-level representations of the interview audio file to produce a person-level embedding. These embeddings were used to perform 10-fold cross-validation with a ridge regression model, and its performance was measured using Pearson correlation coefficient ($r$) and mean squared error ($mse$).
The success of WhiSPA-394 can be attributed to its integration of psychological feature alignment, which complements semantic alignment by explicitly encoding affective dimensions such as valence and arousal. The improvements in outcomes like INT and DIS further support this interpretation since these constructs often rely on subtle vocal cues, such as pause distribution, pitch variability, and vocal tone as established by prior works Kotov et al. (2024). WhiSPA-394 is better equipped to model due to its expanded embedding space. By injecting dimensions with psychological relevance into the alignment process, the model bridges the gap between the prosodic information in speech and the textual semantics used to train baseline models like WhiSA. This dual alignment likely enhances the model’s ability to capture both the what (semantic content) and the how (affective delivery) of speech, enabling more accurate predictions of psychological scales.
@@
Investigation of the choice of alignment objective towards performance (Table 4) revealed that Noise Contrastive Estimation (NCE) consistently produced a better-aligned model than cosine similarity (CS). NCE outperforms CS Equation 1 across all models, likely because NCE optimizes for discriminative learning, encouraging more separation between positive and negative samples in the embedding space. This enhances the model’s ability to encode nuanced semantic and psychological cues, as reflected in WhiSPA-394’s better alignment and task performance. When comparing WhiSPA-394 and WhiSPA-384, we notice the recurring trend with NCE granting a greater optima during alignment than CS as exemplified in Table 4. However, WhiSPA-384 holds its ground in HiTOP, achieving comparable correlations. This suggests that WhiSPA-394’s architecture may generalize well to diverse datasets but thrives in highly semantic and affective audio contexts like WTC. All results noted in Table 2 and Table 3 were from models trained with NCE.
PCL HiTOP VAL
Model INT DIS THD (segment)
@@
WhiSPA captures semantic features of its text-based teacher without needing additional SBERT representations.
Adding SBERT representations to WhiSPA offers little benefit, as it already captures nearly all the information from its teacher model. The results in Table 5 underscore the slight increase in correlations after adding text-based embeddings from SBERT-384 with WhiSPA, as seen in the last row. WhiSPA, trained through a student-teacher alignment paradigm, appears to achieve a semantic and psychological optimum during convergence. This is evident in its substantial performance gains over Whisper, which lacks the semantic and psychological depth provided by language models.
Notably, WhiSPA-394 demonstrates clear improvements in specific measures such as INT and VAL, with gains of $+3$ and $+6$ Pearson points, respectively, when compared to SBERT-384. While these increments may seem marginal, they substantiate our claim that WhiSPA-394 effectively captures nearly all the information encoded by its text-based teacher model, SBERT-384. Further, when WhiSPA-394 representations are augmented with SBERT embeddings, the downstream task performance exhibits statistically insignificant gains, suggesting minimal benefit.
Ultimately, these findings highlight an important observation: The potential of cross-modal alignment may be constrained by the representational efficacy of the teacher model. The marginal returns from text-based augmentation indicate that WhiSPA has already learned to encode the critical semantic and psychological cues provided by its teacher, reflecting the success of the alignment process.
@@ 6 Conclusion
Our findings demonstrate that WhiSPA effectively integrates semantic and psychological information from speech, enhancing state-of-the-art audio representations for psychological and mental health assessments. By aligning Whisper’s audio embeddings with SBERT’s text embeddings enriched with psychological features, we found not only consistent improvement for ten self-supervised tasks but also significantly greater accuracies over 11 downstream person-level psychological tasks as compared to modern audio models. Finally, we found little benefit in adding SBERT to WhiSPA embeddings. This suggests that the cross-modal alignment objective we employed affords audio-encoder-decoder models like Whisper the advantages of text-only transformer-LMs that have been trained on much larger datasets than audio foundation models. We see this as part of a growing body of work to create holistic multi-modal foundation models that result in a richer and more authentic representation of human communication.
@@ 7 Limitations
Currently, WhiSPA’s training scope idealizes lexically derived psychological features, which can negatively impact its ability to discern acoustic information. Although we have shown drastic improvements in yielding semantically contextual audio embeddings, it begs the question of how much more affectively contextual the embeddings can be for psychological predictions and emotion recognition tasks. This would require injecting psychological features that are not only lexically derived but also acoustically–since acoustics and vocal prosody are known to convey more affect Low et al. (2020).
Subsequently, since the alignment procedure places equal weight on all dimensions of embedding, the semantic dimensions outnumber the psychological dimensions. This is likely resulting in the convergence of WhiSPA on a semantic similarity optima. Potentially, we would like to explore a multi-objective weighted loss criteria where parameters can be tuned for placing higher priority on affect over semantics in some cases.
@@ 8 Ethical Implications
The WTC and HiTOP recordings took place in a clinical setting at the Stony Brook WTC Health and Wellness Program. Each participant gave informed consent and was fully informed about the study, that it was voluntary to take part, and that they had the right to withdraw at any time without giving a reason or that it would affect their treatment. After the interview, participants were debriefed (for more details about the WTC data collection, see Kjell et al. (2024); for more details about the HiTOP data, see Kotov et al. (2022, 2024). The studies and data uses were approved by the Institutional Review Board at an undisclosed university for privacy reasons.
@@ A.3 Annotations
The documentation accompanying the iHiTOP interview dataset was utilized to report the coverage of its domains, demographic information, and other relevant details. The dataset’s focus on structured psychological interviews and its linguistic properties were described in the paper to contextualize its relevance to this research. This information was presented to ensure transparency and reproducibility. The WTC dataset assessed PTSD symptom severity and related constructs, including anxiety and depression, using English-language data from WTC emergency responders in the Stony Brook Health Program. Linguistic features such as RoBERTa-large embeddings, n-grams, and LDA topics were used to analyze behavioural patterns alongside closed-vocabulary features like pronouns and death-related terms (LIWC-22). The development dataset included 1,437 participants, and the prospective dataset included 346, with a mean age of 58 years, predominantly male (93% and 91%, respectively) and white (54% and 49%). The analysis emphasized language markers of stress, anxiety, and trauma while reflecting on participants’ experiences of 9/11. Ethical safeguards, including IRB approval, informed consent, and automated anonymization, ensured compliance. While comprehensive in its linguistic and demographic scope, the study was limited to English speakers and WTC responders, constraining generalizability.
@@
Figure 7: Annotator’s affective circumplex visual grid for the task of manually annotating acoustic segments of speech from both datasets.

### v2

@@ 5 Results & Discussion
We consider three popular, robust speech encoders as baselines: Wav2Vec2-BERT [^4] W2V2B version: wav2vec2-bert-CV16-en  Communication et al. (2023); Chung et al. (2021), HuBERT [^5] HuBERT version: hubert-large-ls960-ft Hsu et al. (2021), and Whisper Radford et al. (2022), which are referred to as W2V2B, HuBERT, and Whisper-384, respectively. We measured the effectiveness of these embeddings by computing Pearson correlation coefficient ($r$) and mean squared error ($mse$) over a 10-fold cross-validated ridge regression model for each task. For each model variant, we encode audio segments for each participant and aggregate them with a statistical mean to represent person-level embeddings for the tasks in Table 2 and Table 3.
@@
Interestingly, deriving psychological estimates from semantic dimensions (WhiSPA-394) was consistently better than the replacement (WhiSPA-384r) of 10 semantic dimensions with PsychEmb dimension. This shows the importance of curating the semantic dimensions before replacing them with different embeddings.
@@
Figure 3: Bivariate KDE contour plot of PCA dimensionally reduced speech/text embeddings. Speech representations in blue. Text representations in red.
@@
Model HiTOP WTC INT DIS ANT SOM THD DET PCL REX AVO NAM HYP $r(\uparrow)$ $mse(\downarrow)$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ $r$ $mse$ W2V2B .50 .17 .46 .21 .35 .11 -.00 .24 .27 .11 .32 .20 .14 133.19 .14 12.08 .07 3.99 .13 10.98 .10 17.17 HuBERT .50 .17 .53 .19 .36 .11 .07 .23 .28 .11 .31 .20 .21 129.86 .22 11.72 .07 3.99 .19 10.80 .15 16.99 Whisper-384 .39 .19 .33 .24 .33 .11 .07 .23 .28 .11 .29 .20 .23 128.85 .21 11.77 .06 4.00 .19 10.87 .23 16.41 WhiSA-384 .55$\dagger$ .16 .53$\dagger$ .19 .43$\dagger$ .10 .22$\dagger$ .23 .37$\dagger$ .10 .33$\dagger$ .18 .29$\dagger$ 119.68 .27$\dagger$ 11.26 .19$\dagger$ 3.90 .26$\dagger$ 10.12 .28$\dagger$ 15.56 WhiSPA-384r .56$\dagger$ .15 .53$\dagger$ .19 .42$\dagger$ .10 .23* .22 .39$\dagger$ .10 .39$\dagger$ .19 .34$\dagger$ 119.24 .30$\dagger$ 11.23 .17 3.88 .31$\dagger$ 10.08 .32$\dagger$ 15.54 WhiSPA-394 .57$\dagger$ .15 .54$\dagger$ .19 .43$\dagger$ .10 .$\dagger .2$37$\dagger$ .10 .38$\dagger$ .19 .35$\dagger$ 118.91 .30$\dagger$ 11.18 .20 3.85 .32$\dagger$ 10.09 .32$\dagger$ 15.48
[…]
@@
Semantic-Psychological alignment is SotA for audio-based psychological assessments.
Table 3 shows that the improvements brought by our aligned models over traditional models were preserved even when evaluated on a spectrum of downstream psychological assessment tasks. In particular, the alignment showed a stark increase in capturing deeper psychological conditions such as INT (internalizing) ($\geq 16$ Pearson points) and DIS (disinhibition) ($\geq 20$ Pearson points) from very long durations of speech data. Consistent with behaviours exhibited with PsychEmb dimensions, in Table 2, semantic-psychological alignment from semantically-derived psychological dimensions (WhiSPA-394) performed the best, followed by semantic-psychological alignment from replacement (WhiSPA-384r) and finally semantic-only alignment (WhiSA-384). For these tasks, we averaged the segment-level representations of the interview audio file to produce a person-level embedding. These embeddings were used to perform 10-fold cross-validation with a ridge regression model, and its performance was measured using Pearson correlation coefficient ($r$) and mean squared error ($mse$).
The success of WhiSPA-394 can be attributed to its integration of psychological feature alignment, which complements semantic alignment by explicitly encoding affective dimensions such as valence and arousal. The improvements in outcomes like INT and DIS further support this interpretation since these constructs often rely on subtle vocal cues, such as pause distribution, pitch variability, and vocal tone as established by prior works Kotov et al. (2024). By injecting dimensions with psychological relevance into the alignment process, the model bridges the gap between the prosodic information in speech and the textual semantics used to train baseline models like WhiSA. This dual alignment likely enhances the model’s ability to capture both the what (semantic content) and the how (affective delivery) of speech, enabling more accurate predictions of psychological scales.
@@
Investigation of the choice of alignment objective towards performance (Table 4) revealed that Noise Contrastive Estimation (NCE) consistently produced a better-aligned model than cosine similarity (CS). This is likely because NCE optimizes for discriminative learning, encouraging more separation between positive and negative samples in the embedding space Ye et al. (2022), enhancing the model’s ability to encode nuanced semantic and psychological cues. When comparing WhiSPA-394 and WhiSPA-384, we notice the recurring trend with NCE granting a greater optima during alignment than CS as exemplified in Table 4. However, WhiSPA-384 holds its ground in HiTOP, achieving comparable correlations. This suggests that WhiSPA-394’s architecture may generalize well to diverse datasets but thrives in highly semantic and affective audio contexts like WTC.
Model PCL HiTOP VAL
INT DIS THD (segment)
@@
WhiSPA captures semantics without the need for appending SBERT representations.
The last row in Table 5 underscores the marginal increase in correlations after appending SBERT embeddings to WhiSPA. WhiSPA, trained through a student-teacher alignment paradigm, appears to reach a semantic and psychological optimum during convergence. This is evident in its substantial performance gains over Whisper, which lacks the semantic and psychological depth provided by language models. However, the potential of cross-modal alignment may be constrained by the representational efficacy of the teacher model(s). On human-annotated audio segments, all of the WhiSPA variants achieve substantial improvements in capturing acoustic valence. In comparison with Whisper-384, WhiSPA-384r exhibits a gain of $+15$ Pearson points in VAL (acoustic valence) which exemplifies the reduction in the semantic/psychological gap between audio models and text-based models. Notably in Figure 8, WhiSPA-394 demonstrates clear improvements in specific measures such as INT and VAL, with gains of $+3$ and $+8$ Pearson points, respectively, when compared to its teacher, SBERT-384.
Ultimately, these findings highlight two important observations: (1) WhiSPA effectively captures nearly all the information encoded by its text-based teacher model, SBERT. (2) The marginal returns from appending text-based representations indicate that WhiSPA successfully learns to encode the critical semantic and psychological cues provided by its teachers, reflecting the success of the distillation.
WhiSPA’s representations are interpretable through language semantically associated with psychological dimensions.
6(a) shows that n-grams known to be indicative of PTSD severity from prior studies Kjell et al. (2024) —- including first-person pronouns, experienced symptoms, psychological distress, and negative affect – yield significantly higher correlations with WhiSPA’s predictions compared to Whisper. In contrast, 6(b) reveals that language discussing relationships and positive affect is more negatively associated with WhiSPA’s scores. These findings indicate that the contrastive loss training effectively aligns the latent space with rich semantic and psychological representations, capturing psychologically relevant linguistic markers more robustly. The highly semantic latent spaces of text-based LMs are reflected in WhiSPA’s representations, especially for psychological nuances in spoken language. More quantitative analysis of our model can be found in Appendix subsection A.4
@@ 6 Conclusion
We claim that WhiSPA is a significant step toward more accurate representations of human communication by addressing the modal gap between text and audio, as language models often outperform audio models in predicting psychological attributes. By aligning WhiSPA’s representations with SBERT’s representations enriched with PsychEmb, we found consistent improvement for ten self-supervised tasks and significantly greater accuracies over 11 downstream psychological tasks. We observed only marginal improvements when appending SBERT representations to WhiSPA’s, implying that the distillation process effectively captures the semantic features provided by the teacher language model. Our findings exemplify WhiSPA’s effectiveness in extracting semantic and psychological features from speech, enhancing SotA audio representations for psychological and mental health assessments.
@@ 7 Limitations
While WhiSPA demonstrates significant advancements in providing semantically enriched audio embeddings, its current training paradigm predominantly aligns with psychological features derived from text, potentially limiting its capacity to capture critical acoustic information. This lexical bias, while beneficial for aligning with language-based models, raises an important question: to what extent can WhiSPA’s embeddings be further refined to incorporate affective context for psychological prediction? Given that vocal prosody and acoustic features convey essential emotional and psychological cues beyond textual content Low et al. (2020), incorporating these dimensions is crucial for a more comprehensive representation.
We acknowledge that this strong alignment with text-based language models may introduce an imbalance, diminishing the richness of acoustic cues that are particularly valuable for affective and psychological assessments. Despite WhiSPA’s demonstrated success-—matching its language model teacher in psychological prediction and surpassing state-of-the-art audio models—-there remains an opportunity to enhance its representational capacity by preserving acoustic features. To address this, future work will explore a multi-weighted dual loss objective, ensuring that WhiSPA retains a broader spectrum of information beyond textual representations. We suspect this refinement would not only improve its efficacy in psychological modeling but also enhance its versatility for general-purpose speech tasks like automatic speech recognition (ASR) and emotion recognition in conversation (ERC), where both linguistic and acoustic cues are essential.
@@ 8 Ethical Implications
The WTC and HiTOP recordings took place in a clinical setting where each participant gave consent and was fully informed about the study, that it was voluntary to take part, and that they had the right to withdraw at any time without giving a reason or that it would affect their treatment. After the interview, participants were debriefed (for more details about the WTC data collection, see Kjell et al. (2024); for more details about the HiTOP data, see Kotov et al. (2022, 2024). The studies and data uses were approved by the Institutional Review Board at an undisclosed university for privacy reasons.
@@ A.3 Annotations
The documentation accompanying the iHiTOP interview dataset was utilized to report the coverage of its domains, demographic information, and other relevant details. The dataset’s focus on structured psychological interviews and its linguistic properties were described in the paper to contextualize its relevance to this research. This information was presented to ensure transparency and reproducibility. The WTC dataset assessed PTSD symptom severity and related constructs, including anxiety and depression, using English-language data from WTC emergency responders. Linguistic features such as RoBERTa-large embeddings, n-grams, and LDA topics were used to analyze behavioural patterns alongside closed-vocabulary features like pronouns and death-related terms (LIWC-22). The development dataset included 1,437 participants, and the prospective dataset included 346, with a mean age of 58 years, predominantly male (93% and 91%, respectively) and white (54% and 49%). The analysis emphasized language markers of stress, anxiety, and trauma while reflecting on participants’ experiences of 9/11. Ethical safeguards, including IRB approval, informed consent, and automated anonymization, ensured compliance. While comprehensive in its linguistic and demographic scope, the study was limited to English speakers and WTC responders, constraining generalizability.
@@
Figure 7: Annotator’s affective circumplex visual grid for the task of manually annotating speech segments.
A.4 Quantitative Analysis
PsychEmb’s lower correlations in Figure 8 should not be mistaken for poor performance. With only 10 dimensions, PsychEmb representations achieve a staggering $24$ and $22$ Pearson points on INT and DIS respectively, emphasizing its validity as the psychological teacher. WhiSPA’s consistent improvement over the audio models is attributed to the semantic and psychological dimensions that SBERT and PsychEmb offer. Notably, WhiSPA exemplifies drastic improvements in prediction accuracy for VAL, INT, THT, and PCL compared to Whisper-384.
While WhiSPA demonstrates substantial advancements, surpassing even its text-based LM teacher, SBERT-384, it remains inherently constrained by the representational capacity of the teacher model. If the teacher’s capabilities are limited, these deficiencies inevitably carry over to the student, even after distillation. This is evident in the ARO column, where arousal —- an affective dimension -— is more accurately conveyed through acoustic cues. However, WhiSPA struggles to capture and preserve the acoustic information, instead predominantly aligning with the semantic representations provided by SBERT, thus limiting its ability to fully represent the nuanced affective content inherent in speech.
Beyond demonstrating superior alignment with established PTSD markers, Table 6 highlights WhiSPA’s enhanced sensitivity to psychologically meaningful language patterns. 6(a) shows that n-grams reflecting personal experiences, self-referential content (e.g., first-person pronouns), and negative affective states correlate more strongly with WhiSPA’s predictions than with those of Whisper. WhiSPA appears better attuned to indicators of psychological distress, anxiety, and trauma symptoms—-an advantage likely stemming from the contrastive alignment objective with text-based representations. The model’s capacity to detect nuanced emotional and cognitive expressions in spoken language is further supported by its higher effect sizes on known PTSD-relevant n-grams, underscoring that semantically oriented embeddings can bolster the recognition of clinically significant markers in audio data.
Meanwhile, 6(b) points to a distinctive negative association between WhiSPA’s predicted severity scores and n-grams referencing positive affect or social relationships. This result suggests that the same semantically focused latent space that amplifies negative or distress-related terms also filters out language tied to more adaptive or supportive experiences. In practical terms, such an effect could be advantageous for screening or early detection: positive affect or relational talk might serve as a buffer or resilience indicator, thereby inversely correlating with predicted symptom severity. Taken together, these findings highlight the unique strength of WhiSPAs in capturing a wide spectrum of psychologically relevant linguistic markers, surpassing the granularity offered by audio models alone.
Figure 8: WhiSPA Closes the Semantic/Psychological Representation Gap. WhiSPA consistently outperforms every baseline audio model and, in some cases, even exceeds the performance of the text-based language model teacher.
n-gram r (WhiSPA) r (Whisper)
me 0.261 0.211
ptsd 0.226 0.126
mental 0.200 0.076
because 0.195 0.190
therapist 0.188 0.088
anxiety 0.187 0.075
my therapist 0.175 0.089
my mental health 0.167 0.072
stress 0.165 0.055
want 0.161 0.098
through this 0.160 0.082
pain 0.158 0.171
body 0.156 0.105
this 0.155 0.113
mental health , 0.152 0.051
i had no 0.151 0.135
depression 0.148 0.101
shit 0.147 0.148
but i can’t 0.145 0.041
flashbacks 0.144 0.113
n-gram r (WhiSPA) r (Whisper)
family -0.264 -0.200
will be -0.201 -0.108
college -0.199 -0.099
we’ve -0.190 -0.155
will -0.182 -0.065
wife -0.180 -0.068
pretty -0.176 -0.161
as -0.172 -0.127
good -0.170 -0.170
hopefully -0.167 -0.155
my wife -0.165 -0.070
graduated from -0.163 -0.028
would -0.159 -0.117
able -0.154 -0.051
i would say -0.153 -0.098
able to -0.153 -0.054
kids will -0.153 -0.102
would say -0.152 -0.100
vacations -0.151 -0.152
lucky -0.150 -0.101
Table 6: (a) Top positively correlated N-grams with WhiSPA prediction for PCL scores on the WTC dataset and the corresponding correlations with Whisper predictions. (b) Top negatively correlated N-grams with WhiSPA prediction for PCL scores on the WTC dataset and the corresponding correlations with Whisper predictions. All correlations are statistically significant (p<.05; Benjamini Hochberg corrected).